Please add the following sentence to the question: The company just paid $5 dividend. Question #5: SAPS company stocks are
Answer #1

As per Dividend Discount Model,

Current price of the share P0= D1/r-g Where D1= Dividend in year 1, r= Required rate of return and g= constant growth rate in dividend.

Therefore, D1= D0*g where D0= Current (Last year) Dividend

Hence, P0= [D0*(1+g)]/(r-g)

P0*r – P0*g = D0 + D0*g

P0*r – D0= P0*g – D0*g = g*(P0-D0)

Therefore, g= (P0*r –D0)/ (P0-D0)

Given, P0= $25 and r= 20% and D0=$5

Substituting these values,

Constant Growth rate (g) = (25*0.20-$5)/25-5 = 0/20 = 0